# SAMBO-RL: Shifts-aware Model-based Offline Reinforcement Learning

### v2

@@ Abstract
Model-based offline reinforcement learning trains policies using pre-collected datasets and learned environment models, eliminating the need for direct real-world environment interaction. However, this paradigm is inherently challenged by distribution shift (DS). Existing methods address this issue by leveraging off-policy mechanisms and estimating model uncertainty, but they often result in inconsistent objectives and lack a unified theoretical foundation. This paper offers a comprehensive analysis that disentangles the problem into two fundamental components: model bias and policy shift. Our theoretical and empirical investigations reveal how these factors distort value estimation and restrict policy optimization. To tackle these challenges, we derive a novel Shifts-aware Reward (SAR) through a unified probabilistic inference framework, which modifies the vanilla reward to refine value learning and facilitate policy training. Building on this, we introduce Shifts-aware Model-based Offline Reinforcement Learning (SAMBO-RL), a practical framework that efficiently trains classifiers to approximate SAR for policy optimization. Empirical experiments show that SAR effectively mitigates DS, and SAMBO-RL achieves superior or comparable performance across various benchmarks, underscoring its effectiveness and validating our theoretical analysis.
Machine Learning, ICML
1 Introduction
@@
Offline reinforcement learning (RL) (Lange et al., 2012; Levine et al., 2020) learns policies from the offline dataset generated by a behavior policy, avoiding additional online interaction with the environment. This approach shows remarkable potential in data-driven decision scenarios (Emerson et al., 2023; Sinha et al., 2022) where exploration is costly or hazardous. The model-based framework (Luo et al., 2018) is particularly effective for offline RL (Yu et al., 2020, 2021; Sun et al., 2023; Li et al., 2024). It involves learning models of the environment from the offline dataset and utilizing these models to generate data for policy training. However, directly learning a policy using offline and synthetic data introduces the distribution shift challenge. This shift causes the training objective to deviate from the true objective of RL, leading to poor test performance.
@@
Prior methods in model-based offline RL focus on directly leveraging off-policy methods in online RL and heuristic model uncertainty for conservative learning to mitigate the impact of distribution shift. For instance, MOPO (Yu et al., 2020) uses the predicted variance of the learned model as an uncertainty penalty on rewards. MOBILE (Sun et al., 2023) penalizes value learning of model data according to the uncertainty of the Bellman Q-function estimated by ensemble models. RAMBO (Rigter et al., 2022) considers the trajectory distribution in the worst case and adversarially trains the policy and transition model while ensuring accurate predictions. However, these approaches have resulted in a biased training objective due to the model uncertainty and remain the distribution shift problem to be solved.
@@
Figure 1: The negative impacts of model bias and policy shift. (a) In the 1D grid world, the agent can take two actions: move left (L) or move right (R) from the current state $S$ to the adjacent grid cells $S-1$ or $S+1$, with movement obstructed at the boundaries. The initial state is uniformly distributed, and the agent receives undiscounted rewards upon reaching specific states. (b) The left panel illustrates value estimates during training and the final policy convergence with and without model bias. Even slight model bias can significantly distort value function estimates, leading the learned policies to deviate from the true optimal policy in the actual environment. (c) The right panel showcases expected returns and the KL divergence between the learned policy $\pi$ and the behavior policy $\pi_{b}$ during training with and without policy shift. We find that policy shift implicitly restricts $\pi$ to remain close to $\pi_{b}$, thereby slowing policy convergence.
@@
In this paper, we demystify the distribution shift problem in model-based RL and study its fundamental properties. We first show that this problem arises from both model bias and policy shift, and further analyze their impacts on policy optimization. Unlike the heuristic descriptions of model uncertainty, model bias fundamentally characterizes the inconsistency between the learned model and actual environment dynamics in predicting subsequent states. This bias leads to inaccurate estimations of trajectory returns, resulting in distorted objectives. If the learned model overestimates or underestimates the probability of generating a trajectory, it consequently causes the overestimation or underestimation of trajectory returns. Additionally, policy shift depicts the changes in state-action visitation distribution caused by the deviation between the learned policy $\pi$ and behavior policy $\pi_{b}$. This shift implicitly forces $\pi$ towards $\pi_{b}$, hindering the learning process and slowing down policy convergence. In contrast, previous studies heuristically mitigate these two issues from off-policy algorithms and model uncertainty.
@@
To address these challenges, we theoretically derive adjustments for model bias and policy shift by employing a unified probabilistic inference framework, and introduce a shifts-aware reward, which integrates the vanilla reward with adjustments for model bias and policy shift. The model bias adjustment penalizes the reward to mitigate issues of overestimation or underestimation, thereby refining value learning and preventing its convergence to local optima. Meanwhile, the policy shift modification revises the reward to motivate $\pi$ to escape the limitation of $\pi_{b}$, accelerating policy convergence. Building upon our theoretical framework, we develop Shifts-aware Model-based Offline Reinforcement Learning (SAMBO-RL). This approach incorporates model bias adjustment for model-generated data and policy shift modification for offline data. Furthermore, it learns a transition classifier and action classifiers to approximate the shifts-aware reward and then performs policy optimization based on this adjusted reward. Our contributions include:
@@
  • We demystify the distribution shift problem in model-based offline RL, i.e., model bias distorts the estimation of the value function, while policy shift imposes an implicit policy limitation that hinders policy convergence.
@@
  • We upgrade the vanilla reward into a shifts-aware reward incorporating model bias adjustment and policy shift modification. This modified reward is used for policy learning to mitigate distribution shift in model-based offline RL.
@@
  • We devise SAMBO-RL, which trains transition and action classifiers to approximate the shifts-aware reward, effectively addressing both model bias and policy shift.
@@
  • We show that shifts-aware reward effectively mitigates distribution shift, and SAMBO-RL demonstrates superior or comparable performance across various benchmarks.
2 Related Work
Model-free Offline RL encompasses two primary approaches: policy constraints and value regularization, both incorporating conservatism (Jin et al., 2021). Policy constraint methods aim to keep the learned policy close to the behavioral policy, mitigating distribution shift. For instance, BEAR (Kumar et al., 2019) and BRAC (Wu et al., 2019) constrain the learned policy by minimizing different divergences. BCQ (Fujimoto et al., 2019) perturbs actions during learning to ensure they remain within the action space of the offline dataset by learning a generative model of the behavior policy. On the other hand, value regularization methods add conservative regularization terms to the value optimization objective, avoiding out-of-distribution (OOD) actions. CQL (Kumar et al., 2020) penalizes the Q-values of OOD samples by incorporating regularization into the value loss function. TD3+BC (Fujimoto & Gu, 2021) introduces a behavior cloning regularization into the TD3 (Fujimoto et al., 2018) objective. Similarly, EDAC (An et al., 2021) and PBRL (Bai et al., 2021) penalize the Q-function based on uncertainty derived from an ensemble of Q-networks. MCQ (Lyu et al., 2022) actively trains OOD actions by constructing pseudo-target values to alleviate over-pessimism.
Model-based Offline RL approximates the environment using learned models and performs conservative policy optimization (Lu et al., 2021), achieving high data efficiency (Li et al., 2024). MOPO (Yu et al., 2020) incorporates conservatism by penalizing rewards based on the uncertainty of model predictions. MOReL (Kidambi et al., 2020) introduces a pessimistic MDP to penalize rewards of state-action pairs in unexplored regions. COMBO (Yu et al., 2021) extends CQL to the model-based setting, regularizing the value function on OOD samples. TT (Janner et al., 2021) uses a transformer to model offline trajectories and employs beam search for planning. RAMBO (Rigter et al., 2022) adopts an adversarial approach, training policy and model within a robust framework while ensuring accurate transition predictions of the model. CBOP (Jeong et al., 2022) adaptively weights multi-step returns in a model-based value expansion framework (Feinberg et al., 2018). Lastly, MOBILE (Sun et al., 2023) introduces penalties during value learning by leveraging the uncertainty of the Bellman Q-function estimates derived from ensemble learned dynamic models.
3 Preliminaries
@@
MDP. We focus on a Markov decision process (MDP) specified by the tuple $M=(\mathcal{S},\mathcal{A},p,r,\mu_{0},\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, $p(s^{\prime}|s,a)$ is the environment transition dynamics, positive $r(s,a)>0$ is the reward function, $\mu_{0}(s)$ is the initial state distribution, and $\gamma\in(0,1)$ is the discount factor. Given an MDP, we define the state value function $V^{\pi}(s)=\mathbb{E}_{\pi,p}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s\right]$ and the $Q$ function $Q^{\pi}(s,a)=\mathbb{E}_{\pi,p}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a\right]$.
@@
MBPO. Similar to previous works (Yu et al., 2020, 2021; Rigter et al., 2022; Sun et al., 2023), we focus on model-based policy optimization (MBPO) (Janner et al., 2019), which employs an actor-critic RL algorithm. MBPO first performs $k$-step rollouts using the model $m(s^{\prime}|s,a)$ starting from state $s\in\mathcal{D}_{env}$ and adds the generated data to a replay buffer $\mathcal{D}_{m}$. Then it optimizes the policy using mini-batches of data sampled from $\mathcal{D}_{env}\cup\mathcal{D}_{m}$, where each datapoint is sampled from the real dataset $\mathcal{D}_{env}$ with probability $f$ and the model dataset $\mathcal{D}_{m}$ with probability $1-f$.
Trajectory Formulation. To demystify the distribution shift, We consider the trajectory formulation of optimization objectives. Let $p^{\pi}(\cdot)$ represent the probability distribution over trajectories =$\mathcal{H}=\{\tau=(s_{0},a_{0},r_{0},s_{1},...)\}$ generated by executing policy $\pi$ under dynamics $p$. This is given by:
@@
The training data $\mathcal{D}_{train}$ can be interpreted as being generated by a data collection policy $\pi_{c}$ through rollouts in dynamics $q$. Consequently, the probability of observing trajectory $\tau$ in the training data can be formulated as $q^{\pi_{c}}(\tau)$. The reinforcement learning objective and the practical training objective are then formalized as the following equations:
@@
where $R(\tau)=\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})$ represents the discounted return of the whole trajectory $\tau=(s_{0},a_{0},r_{0},s_{1},...)$.
4 Shifts-aware Policy Optimization
@@
We first formalize the distribution shift problem as shift weighting and further delve into its negative effects on policy optimization from theory and experiments. To address these challenges, we employ probabilistic inference to derive a surrogate objective and formulate a shifts-aware reward. Finally, we implement this theoretical framework within model-based offline RL and introduce SAMBO-RL, a practical algorithm designed to mitigate distribution shifts.
4.1 Demystifying Distribution Shift
@@
We examine the distribution shift problem from an optimization perspective and identify its origins as model bias and policy shift. Then we explore how these factors affect accuracy in value estimation and impose implicit policy restrictions, resulting in a biased objective and slow convergence. The supplementary analysis is in Appendix A.1.
@@
Shift Weighting. The training objective $\mathcal{J}(\pi)$ in (2) can be expressed as the following trajectory formulation:
@@
Compared to (2), it indicates that for a given trajectory $\tau$, the true objective under practical sampling distributions differs in shift weighting $\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}$ from the training objective, which formalizes the distribution shift in trajectory formulation.
Substituting (1) into the shift weighting $\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}$, we can attribute this weighting to two primary components: model bias and policy shift, reformulated as the following:
@@
Model bias illustrates the discrepancy between the dynamics used for training and those of the actual environment, while policy shift represents the divergence between the policies used for sampling training data and those being learned.
@@
(i) To illustrate the effect of model bias, we examine the scenario where the training policy $\pi$ operates within model dynamics $q$, which diverges from the actual environment dynamics $p$. In this context, the training objective $\mathcal{J}(\pi)$ is:
@@
$\mathcal{J}(\pi)$ $=\mathbb{E}_{\tau\sim p^{\pi}}\left[R(\tau)\prod\limits_{t=0}^{\infty}\frac{q(s_{t+1}|s_{t},a_{t})}{p(s_{t+1}|s_{t},a_{t})}\right].$
@@
Compared to the true objective (3), the compounded prediction bias of the model significantly affects value estimation and leads to a distorted objective. Specifically, when the model $q$ overestimates or underestimates the probability of generating trajectory $\tau$, it consequently overestimates or underestimates the expected return of this trajectory in the real environment. This essentially introduces bias for the optimization objective, resulting in wrong training directions.
@@
(ii) To delve deeper into the policy optimization process under policy shift, we consider the setting where a sampling policy $\pi_{c}$ is executed within the true environment transition dynamics $p$. In this scenario, the training objective $\mathcal{J}(\pi)$ is:
$\mathcal{J}(\pi)$ $=\mathbb{E}_{(s,a)\sim d^{p,\pi}}\left[r(s,a)\frac{d^{p,\pi_{c}}(s,a)}{d^{p,\pi}(s,a)}\right].$
@@
Here, $d^{p,\pi}(s,a)$ is the state-action visitation distribution under dynamics $p$ and policy $\pi$, i.e. $d^{p,\pi}(s,a)=(1-\gamma)\mathbb{E}_{s_{0}\sim\mu_{0},s_{t}\sim p,a_{t}\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{I}(s_{t}=s,a_{t}=a)\right]$, where $\mathbb{I}$ represents the indicator function. Note that $\pi_{c}$ is related to $\pi$. Compared to the true objective, the policy shift implicitly forces policy $\pi$ to approach $\pi_{c}$, hindering the learning process and slowing down policy convergence. Specifically, for a state-action pair $(s,a)$ associated with a high reward, ideal policy updates should increase $\pi(a|s)$. However, if the currently learned policy $\pi$ adequately approximates the optimal policy and satisfies $d^{p,\pi}(s,a)>d^{p,\pi_{c}}(s,a)$, the reward for $(s,a)$ can be underestimated. This situation could even prompt policy updates to decrease $\pi(a|s)$, contrary to the optimal direction. Similarly, for the state-action pair $(s,a)$ with a low reward, if the currently $\pi$ is reasonably good and leads to $d^{p,\pi}(s,a)<d^{p,\pi_{c}}(s,a)$, the reward can be overestimated, potentially resulting in policy updates to increase $\pi(a|s)$ that do not align with optimal policy directions.
@@
Toy Examples. To empirically validate our analysis of model bias and policy shift, we use a 1D grid world example, as depicted in Fig. 1 (a). In Fig. 1 (b), we compare the training process with and without model bias and find the value function overestimation and underestimation under model bias. These results reveal that even slight deviations in model dynamics can significantly distort value function estimates, leading to a biased training objective. Consequently, the policies learned by the agents deviate from the optimal policy. To assess the impact of the policy shift, we examine a model-free offline setting where data is pre-collected using a behavior policy $\pi_{b}$. Fig. 1 (c) showcases that the KL divergence between the learned policy $\pi$ and $\pi_{b}$ under policy shift is substantially reduced compared to the scenario without policy shift, sometimes even decreasing undesirably. This indicates that policy shift does compel $\pi$ to align closely with $\pi_{b}$, thereby impeding the training process and substantially slowing down the policy convergence.
4.2 Shifts-aware Reward via Probabilistic Inference

### v3

# Mitigating Distribution Shift in Model-based Offline RL via Shifts-aware Reward Learning

@@ Abstract.
Model-based offline reinforcement learning trains policies using pre-collected datasets and learned environment models, eliminating the need for direct real-world environment interaction. However, this paradigm is inherently challenged by distribution shift (DS). Existing methods address this issue by leveraging off-policy mechanisms and estimating model uncertainty, but they often result in inconsistent objectives and lack a unified theoretical foundation. This paper offers a comprehensive analysis that disentangles the problem into two fundamental components: model bias and policy shift. Our theoretical and empirical investigations reveal how these factors distort value estimation and restrict policy optimization. To tackle these challenges, we derive a novel shifts-aware reward through a unified probabilistic inference framework, which modifies the vanilla reward to refine value learning and facilitate policy training. Building on this, we develop a practical implementation that leverages classifier-based techniques to approximate the adjusted reward for effective policy optimization. Empirical results across multiple benchmarks demonstrate that the proposed approach mitigates distribution shift and achieves superior or comparable performance, validating our theoretical insights.
Key words and phrases:
Offline RL, Distribution Shift, Model-based RL
1. Introduction
@@
Offline reinforcement learning (RL) Lange et al. (2012); Levine et al. (2020) learns policies from the offline dataset generated by a behavior policy, avoiding additional online interaction with the environment. This approach shows remarkable potential in data-driven decision scenarios Emerson et al. (2023); Sinha et al. (2022) where exploration is costly or hazardous. The model-based framework Luo et al. (2018) is particularly effective for offline RL Yu et al. (2020, 2021); Sun et al. (2023); Li et al. (2024). It involves learning models of the environment from the offline dataset and utilizing these models to generate data for policy training. However, directly learning a policy using offline and synthetic data introduces the distribution shift challenge. This shift causes the training objective to significantly deviate from the true objective of reinforcement learning, leading to poor performance during testing.
@@
Prior methods in model-based offline RL focus on directly leveraging off-policy methods in online RL and heuristic model uncertainty for conservative learning to mitigate the impact of distribution shift. For instance, MOPO Yu et al. (2020) uses the predicted variance of the learned model as an uncertainty penalty on rewards. MOBILE Sun et al. (2023) penalizes value learning of model data according to the uncertainty of the Bellman Q-function estimated by ensemble models. RAMBO Rigter et al. (2022) considers the trajectory distribution in the worst case and adversarially trains the policy and model while ensuring accurate predictions. However, these approaches result in a biased objective by only addressing the distribution shift from model inaccuracies, leaving the broader distribution shift unresolved.
@@
Figure 1. The negative impacts of model bias and policy shift. (a) In the 1D grid world, the agent can take two actions: move left (L) or move right (R) from the current state $S$ to the adjacent grid cells S1S-1 or $S+1$, with movement obstructed at the boundaries. The initial state is uniformly distributed, and the agent receives undiscounted rewards upon reaching specific states. (b) The left panel illustrates value estimates during training and the final policy convergence with and without model bias. Even slight model bias can significantly distort value function estimates, leading the learned policies to deviate from the true optimal policy in the actual environment. (c) The right panel showcases expected returns and the KL divergence between the learned policy and the behavior policy with and without policy shift. We find that policy shift implicitly slows policy convergence.
@@
In this paper, we demystify the distribution shift problem in model-based RL and study its fundamental properties. We first show that this problem arises from both model bias and policy shift, and further analyze their impacts on policy optimization. Unlike the heuristic descriptions of model uncertainty, model bias fundamentally characterizes the inconsistency between the learned model and actual environment dynamics in predicting subsequent states. This bias leads to inaccurate estimations of trajectory returns, resulting in distorted objectives. If the learned model overestimates or underestimates the probability of generating a trajectory, the resulting distribution shift during training consequently leads to overestimation or underestimation of trajectory returns. Additionally, policy shift depicts the changes in state-action visitation distribution caused by the deviation between the training data collection policy and the learned policy. This distribution shift between training and deployment implicitly perturbs policy updates and thereby hinders learning. Existing studies heuristically mitigate these issues with off-policy mechanisms and uncertainty penalties, lacking a consistent, theoretically grounded framework.
@@
To address these challenges, we theoretically derive adjustments for model bias and policy shift by employing a unified probabilistic inference framework, and introduce a shifts-aware reward (SAR), which integrates the vanilla reward with adjustments for model bias and policy shift. The model bias adjustment penalizes the reward to mitigate issues of overestimation or underestimation, thereby refining value learning and alleviating distribution shift during training. Meanwhile, the policy shift modification adjusts the reward to mitigate distribution shift between training and deployment, stabilizing the policy updates and accelerating policy convergence. Building on our theoretical framework, we develop Shifts-aware Model-based Offline Reinforcement Learning (SAMBO-RL). This approach incorporates model bias adjustment for model-generated data and policy shift modification for training data. Furthermore, it learns a transition classifier and action classifiers to approximate the shifts-aware reward and then performs policy optimization based on this adjusted reward. Our contributions include:
@@
  • We demystify the distribution shift in model-based offline RL, i.e., model bias causes distribution inconsistency in the training data, ultimately distorting value estimation, while policy shift leads to distribution inconsistency between training and deployment, hindering policy convergence.
@@
  • We upgrade the vanilla reward into a shifts-aware reward that incorporates model bias adjustment and policy shift modification, effectively mitigating distribution shift.
@@
  • We devise SAMBO-RL, which trains a transition classifier and an action classifier to approximate the shifts-aware reward, effectively addressing both model bias and policy shift.
@@
  • We empirically show that shifts-aware reward effectively mitigates distribution shift, and SAMBO-RL demonstrates comparable performance across various benchmarks.
2. Preliminaries
@@
MDP. We focus on a Markov decision process (MDP) specified by the six-elements tuple $M=(\mathcal{S},\mathcal{A},p,r,\mu_{0},\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, $p(s^{\prime}|s,a)$ is the environment transition dynamics, positive $r(s,a)>0$ is the reward function, $\mu_{0}(s)$ is the initial state distribution, and $\gamma\in(0,1)$ is the discount factor typically used for future reward weighting. Given an MDP, we define the state value function of the policy $\pi$: $V^{\pi}(s)=\mathbb{E}_{\pi,p}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s\right]$ and the $Q$ function of the policy $\pi$: $Q^{\pi}(s,a)=\mathbb{E}_{\pi,p}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a\right]$.
@@
MBPO. Similar to previous works Yu et al. (2020, 2021); Rigter et al. (2022); Sun et al. (2023) in offline RL, we focus on model-based policy optimization (MBPO) Janner et al. (2019), which uses an actor-critic RL algorithm. Define $\mathcal{D}_{env}$ as the offline dataset collected by behavior policy $\pi_{b}$ in the real environment, and model dataset $\mathcal{D}_{m}$ as the buffer storing synthetic data. MBPO first performs $h$-step rollouts using the model $m(s^{\prime}|s,a)$ starting from state $s\in\mathcal{D}_{env}$ and adds the generated data to the buffer $\mathcal{D}_{m}$. It then optimizes the policy using data sampled from $\mathcal{D}_{env}\cup\mathcal{D}_{m}$, where each datapoint is drawn from $\mathcal{D}_{env}$ with probability $f$ and $\mathcal{D}_{m}$ with probability 1f1-f. We denote the data used for training as $\mathcal{D}_{train}$.
[…]
@@
The training data $\mathcal{D}_{train}$ can be interpreted as generated by a data collection policy $\pi_{c}$ through rollouts in an unknown dynamics $q$. $\pi_{c}$ is inherently related to $\pi$, as the dataset $\mathcal{D}_{train}$ includes samples generated by the policy $\pi$. Consequently, the probability of observing trajectory $\tau$ in the training data is formulated as $q^{\pi_{c}}(\tau)$. Then the RL objective and the practical training objective are
@@
[…]
3. Related Work
Model-free Offline RL encompasses two primary approaches: policy constraints and value regularization, both incorporating conservatism Jin et al. (2021). Policy constraints keep the learned policy close to the behavior policy, while value regularization adds conservative terms to the value optimization. BEAR Kumar et al. (2019) and BRAC Wu et al. (2019) constrain the learned policy by minimizing different divergences. BCQ Fujimoto et al. (2019) perturbs actions during learning to ensure they remain within the action space of the offline dataset by learning a generative model of the behavior policy. SACS Mao et al. (2024a) simultaneously performs out-of-distribution (OOD) state correction using a model that is not employed for data generation and OOD action suppression in offline RL. CQL Kumar et al. (2020) penalizes the Q-values of OOD samples by incorporating regularization into the value loss function. TD3+BC Fujimoto and Gu (2021) introduces a behavior cloning regularization into the TD3 Fujimoto et al. (2018) objective. Similarly, EDAC An et al. (2021) and MCQ Lyu et al. (2022b) actively trains OOD actions by constructing pseudo-target values to alleviate over-pessimism. DMG Mao et al. (2024b) leverages mild action generalization and generalization propagation to enhance generalization in offline RL. SEABO assigns larger rewards to transitions closer to their nearest expert demonstrations, alleviating offline RL’s heavy reliance on the offline data annotated with reward labels. A2PR Liu et al. (2024) selects high-advantage actions as an imitation learning regularizer to mitigate performance degradation from excessive conservatism.
Model-based Offline RL approximates the environment using learned models and performs conservative policy optimization Lu et al. (2021), achieving high data efficiency Li et al. (2024). MOPO Yu et al. (2020) incorporates conservatism by penalizing rewards based on the uncertainty of model predictions. MOReL Kidambi et al. (2020) introduces a pessimistic MDP to penalize rewards of state-action pairs in unexplored regions. COMBO Yu et al. (2021) extends CQL to the model-based setting, regularizing the value function on OOD samples. RAMBO Rigter et al. (2022) adopts an adversarial approach, training the policy and model within a robust framework while ensuring accurate transition predictions of the model. CBOP Jeong et al. (2022) adaptively weights multi-step returns in a model-based value expansion framework Feinberg et al. (2018). CABI Lyu et al. (2022a) improves the reliability of synthesized data by training bidirectional dynamics models and trusting only the samples where the forward and backward models are consistent. C-LAP Alles et al. (2024) learns a generative model of the joint distribution of observations and actions, and imposes an implicit constraint on the generated actions to enable constrained policy learning. Lastly, MOBILE Sun et al. (2023) introduces penalties during value learning by leveraging the uncertainty of the Bellman Q-function estimates derived from ensemble models, achieving SOTA performance among these model-based and model-free approaches.
Cross-domain policy transfer aims to transfer a policy learned in a source environment to a target environment with different dynamics. This is similar to model-based approaches, which learn a model to approximate the environment dynamics and derive high-performing policies in the real environment by training within the learned model. GARAT Desai et al. (2020) learns an adversarial imitation-from-observation policy by discriminating between generated actions and target environment actions. DARC Eysenbach et al. (2020) addresses cross-domain RL via reward correction to enable domain adaptation under changing environment dynamics. SAIL Liu et al. (2019) performs state alignment in cross-domain imitation learning to align state sequences across different domains. Similarly, SRPO Xue et al. (2023) introduces a regularization term based on the KL divergence between state distributions of different domains in the RL objective to improve data efficiency. OTDF Lyu et al. (2025) uses cross-domain offline datasets and mitigates dynamics mismatches via optimal transport, facilitating policy learning.
4. Shifts-aware Policy Optimization
@@
We first formalize the distribution shift problem as shift weighting and further delve into its negative effects on policy optimization from theory and experiments. To address these challenges, we employ probabilistic inference to derive a surrogate objective and formulate a shifts-aware reward. Then we implement this theoretical framework within model-based offline RL and introduce SAMBO-RL, a practical algorithm to mitigate distribution shifts. Finally, we provide a theoretical analysis of SAR.
4.1. Demystifying Distribution Shift
@@
We examine the distribution shift problem from an optimization perspective and identify its origins as model bias and policy shift. Then we explore how these factors affect accuracy in value estimation and induce instability in policy updates.
@@
Shift Weighting. The training objective $\mathcal{J}(\pi)$ in (2) can be expressed as the trajectory formulation:
@@
Compared to $\mathcal{J}_{M}(\pi)$ in (2), it indicates that for a given trajectory $\tau$, the true objective under practical sampling distributions differs in shift weighting $\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}$ from the training objective, which formalizes the distribution shift in trajectory formulation. Substituting (1) into the shift weighting $\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}$, we can attribute it to two primary components: model bias and policy shift. Then the shift weighting can be reformulated as the following:
@@
Model bias illustrates the discrepancy between the dynamics used for training and those of the actual environment, while policy shift represents the divergence between the policies used for sampling training data and those being learned. Both components ultimately lead to the distribution shift issue.
@@
(i) To illustrate the effect of model bias, we examine the synthetic data generation scenario in model-based approaches, where the training policy $\pi$ operates within model dynamics $m$ (i.e. $q=m$), which diverges from the environment dynamics $p$. In this context, according to (4), the training objective $\mathcal{J}(\pi)$ is:
@@
$\mathcal{J}(\pi)$ $=\mathbb{E}_{\tau\sim p^{\pi}}\left[R(\tau)\prod\limits_{t=0}^{\infty}\frac{m(s_{t+1}|s_{t},a_{t})}{p(s_{t+1}|s_{t},a_{t})}\right].$
@@
Compared to the true objective, the compounded prediction bias of the model significantly affects value estimation and leads to a distorted objective. Specifically, when the model $m$ overestimates or underestimates the probability of generating trajectory $\tau$, it consequently overestimates or underestimates the expected return of this trajectory in the real environment. The inconsistency between the model and the environment in predicting future dynamics leads to a mismatch in trajectory distributions during training. This distribution shift in training introduces bias into the optimization objective, thereby guiding the policy toward suboptimal or incorrect updates.
@@
(ii) To delve deeper into the policy optimization process under policy shift, We consider a typical offline reinforcement learning scenario, in which policy optimization is constrained to the training dataset $\mathcal{D}_{train}$, and assume that the learned model perfectly fits the environment dynamics. In this scenario, the training objective $\mathcal{J}(\pi)$ can be formulated as:
$\mathcal{J}(\pi)$ $=\mathbb{E}_{(s,a)\sim d^{p,\pi}}\left[r(s,a)(d^{p,\pi_{c}}(s,a)/d^{p,\pi}(s,a))\right].$
Here, $d^{p,\pi}(s,a)$ is the state-action visitation distribution under dynamics $p$ and policy $\pi$
$d^{p,\pi}(s,a)=(1-\gamma)\mathbb{E}_{s_{0}\sim\mu_{0},s_{t}\sim p,a_{t}\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{I}(s_{t}=s,a_{t}=a)\right],$
@@
where $\mathbb{I}$ represents the indicator function. Note that $\pi_{c}$ is related to $\pi$. Compared to the true objective, the policy shift implicitly forces policy $\pi$ to approach $\pi_{c}$, perturbing policy updates. Specifically, for a state-action pair $(s,a)$ associated with a high reward, ideal policy updates should increase $\pi(a|s)$. However, if the currently learned policy $\pi$ adequately approximates the optimal policy and satisfies $d^{p,\pi}(s,a)>d^{p,\pi_{c}}(s,a)$, the reward for $(s,a)$ can be underestimated. This situation could even prompt policy updates to decrease $\pi(a|s)$, contrary to the optimal direction. Similarly, for the state-action pair $(s,a)$ with a low reward, if the currently $\pi$ is reasonably good and leads to $d^{p,\pi}(s,a)<d^{p,\pi_{c}}(s,a)$, the reward can be overestimated, potentially resulting in policy updates to increase $\pi(a|s)$ that do not align with optimal policy directions. The mismatch between the deployed policy and the training distribution can disrupt updates.
@@
Toy Examples. To empirically validate our analysis of model bias and policy shift, we use a 1D grid world example depicted in Fig. 1 (a). In Fig. 1 (b), we compare the training process with and without model bias and find the value function overestimation and underestimation under model bias. These results reveal that even slight deviations in model dynamics can significantly distort value function estimates, leading to a biased training objective. Consequently, the policies learned by the agents deviate from the optimal policy. To assess the impact of the policy shift, we examine a model-free offline setting where offline data is pre-collected using a behavior policy $\pi_{b}$. Fig. 1 (c) showcases that the KL divergence between the learned policy $\pi$ and behavior policy $\pi_{b}$ under policy shift is substantially reduced compared to the scenario without policy shift, sometimes even decreasing undesirably. This indicates that the policy shift is impeding the training process and substantially slowing down the policy convergence.
4.2. Shifts-aware Reward via Probabilistic Inference